Determine the molar solubility of CaSO4 in a solution containing 9.0×10−2 M K2SO4. Ksp (CaSO4) = 7.1×10-5.

8.4×10-3 M
5.0×10-9 M
7.9×10−4 M
2.2×10−3 M
6.4×10−6 M
Answer #1

2 - (s) Casoy - Ca 2+ + son 2- (ans (aq) s (5+9.0x102) Kąp= s(s+9.0x10-2) J.1810-5 = 5² + 9.0x10? >> 527910x10s - 7.1x16520